HYDRATION EFFECTS ON ELECTRONIC AND VIBRATIONAL STRUCTURE OF PHENOXYL RADI-CALS AND THEIR PARENT ANIONS

\underline{G. N. R. TRIPATHI}, Radiation Laboratory, University of Notre Dame, Notre Dame, IN 46556.